# 4D Gaussian Splatting: Modeling Dynamic Scenes with Native 4D Primitives

### v1

@@ III-C Training
Optimization
In optimization, we only use the rendering loss as supervision. In most cases, combining the representation introduced above with the default training schedule as in [24] is sufficient to yield satisfactory results. However, for scenes with more dramatic changes, we observe issues such as temporal flickering and jitter. We consider that this may arise from suboptimal sampling techniques. Rather than adopting the prior regularization, we discover that batch sampling in time turns out to be effective, resulting in more seamless and visually pleasing appearance of dynamic contents.
Densification in spacetime
In terms of density control, simply considering the magnitude of spatial position gradients is insufficient to assess under- and over-reconstruction over time. To address this, we incorporate the gradients of $\mu_{t}$ as an additional density control indicator. Furthermore, in regions prone to over-reconstruction, we employ joint spatial and temporal position sampling during Gaussian splitting.
IV Compact 4D Gaussian splatting
IV-A Analysis on memory bottleneck
4DGS can be interpreted as modeling the motion of dynamic scenes using piecewise-linear function, analogous to how 3D Gaussian model geometry. Given sufficient training views, the piecewise-linear function can offer a better fitting capability than those modeling the motion with trigonometric or polynomial functions since enough piecewise-linear functions can also approximate these functions. However, this leads to a large number of primitives to model the motion of the same object, resulting in a considerable redundancy and substantial storage demands. Due to the lack of smoothness priors inherent in the neural networks, its over-strong modeling capability tends to cause overfitting when reconstructing dynamic objects from the under-constraint monocular video.
To mitigate the aforementioned drawbacks, we consider two perspectives: (i) lowering the encoding cost for the attributes of each Gaussian (Section IV-B); (ii) decreasing the total number of primitives (Section IV-C and IV-D).
IV-B Parameter compression
@@
IV-C Insignificant Gaussian removal
@@
IV-D Structured association aware compression
Within our 4DGS representation, there exists natural association between primitives induced by the motion of an object in monocular dynamic scenes. This structural information can be used for model compression, inspired by [39, 40].
Specifically, we introduce a set of anchor points, each parameterized by a group of fixed basic attributes of original 4D Gaussian, along with a context feature $f\in\mathbb{R}^{32}$ and $k$ offsets in time dimension $o_{t}\in\mathbb{R}^{k}$. Each anchor spawns $k$ neural 4D Gaussians during rendering. The offsets of the attributes for these spawned 4D Gaussians are dynamically obtained from MLPs which take the location, context feature, and time offsets of the anchor as input. Once the neural Gaussians are generated, they can be utilized within the differentiable rendering pipeline detailed in Section III.
In optimization, we initiate a warm-up stage for reconstruction of static regions, where the dual quaternions of 4D Gaussians are set to be equivalent by duplicating the parameters of two rotation heads. To encourage motion to be modeled as different neural Gaussians of the same anchor, we incorporate $\mathcal{L}_{\mathrm{sparse}}$ in equation (26) to suppress the motion induced by 4D rotation. Since the offsets are also responsible for learning the motion with time, the anchor growing adopted in the 3D case [39] is no longer suitable. Instead we employ the densification and pruning scheme [24].
V Urban scene 4D Gaussian
We extend 4DGS to urban scene scenarios where boundedness space and limited views of driving videos present unique challenges, along with no 3D bounding boxes and foreground object segmentation. To address these issues, we tailor a set of customizations. We leverage LiDAR point clouds, typically available with corresponding timestamps aggregated across all training frames, to initialize the 4D position of 4D Gaussians. As the distance to the sky is much bigger than the scene’s scale, we employ a learnable cube map, $M_{\mathrm{sky}}$, to model its appearance, which can be queried by view direction $\mathbf{r}$. For each pixel, the cube map is used to consume the remaining transmittance:
@@ V Urban scene 4D Gaussian
$\mathcal{L}_{\mathrm{urban}}=$ $\lambda_{\mathrm{l1}}\mathcal{L}_{\mathrm{l1}}+\lambda_{\mathrm{d-ssim}}\mathcal{L}_{\mathrm{d-ssim}}+\lambda_{\mathrm{sky}}\mathcal{L}_{\mathrm{sky}}+\lambda_{\mathrm{lidar}}\mathcal{L}_{\mathrm{lidar}}$ (30)
$+$ $\lambda_{\mathrm{covt}}\mathcal{L}_{\mathrm{covt}}+\lambda_{\mathrm{sparse}}\mathcal{L}_{\mathrm{sparse}}+\lambda_{\mathrm{rigid}}\mathcal{L}_{\mathrm{rigid}},$
@@
VI Generative 4D Gaussian
We further expand 4DGS for more challenging content generation, such as video-to-4D generation - taking a single-view video as input to generate a 4D object, allowing novel view synthesis from any viewpoint at any timestamp [14]. Existing methods [14, 74, 73, 76, 75] often incorporate customized designs. For fair comparison, we use only anchored images [74] and score distill sampling (SDS) [65] as supervision, independent of 4D representation.
@@ VI Generative 4D Gaussian
$\mathcal{L}=$ $\lambda_{\text{img}}\sum_{i=1}^{N}\|\hat{\mathcal{I}}_{i}^{\text{ref}}-\mathcal{I}_{i}^{\text{ref}}\|_{2}^{2}+\lambda_{\text{img}}\sum_{i=1}^{N}\sum_{j=1}^{M}\|\hat{\mathcal{I}}_{i,j}-\mathcal{I}_{i,j}\|_{2}^{2}$ (32)
@@
where $\lambda_{*}$ denotes the balancing weight, ^$\hat{\mathcal{I}}_{i}^{\text{ref}}$ and ^$\hat{\mathcal{I}}_{i,j}$ are the rendered images under fixed reference camera poses at timestamp $i$, and $\mathcal{I}^{\text{random}}_{i}$ is the rendered image under randomly sampled camera pose at timestamp $i$. We note rendered images ^$\hat{\mathcal{I}}_{i}^{\text{ref}}$, ^$\hat{\mathcal{I}}_{i,j}$ and $\mathcal{I}^{\text{random}}_{i}$ are 4D representation specific.
VII Segment anything in 4D Gaussian
@@
Scale-gated features
We need to obtain the mask scale $s_{\mathbf{M}}$ for a 2D mask $\mathbf{M}$. To that end, $\mathbf{M}$ is first projected into 4D space to get the projected point cloud $\mathcal{P}$. Then $s_{\mathbf{M}}$ is calculated from $\mathcal{P}$ as:
@@
$2\sqrt{\text{std}(\mathcal{X}(\mathcal{P}))^{2}+\text{std}(\mathcal{Y}(\mathcal{P}))^{2}+\text{std}(\mathcal{Z}(\mathcal{P}))^{2}+\text{std}(\mathcal{T}(\mathcal{P}))^{2}}$
@@
The 2D feature $\mathbf{F}^{s}$ at a specific view could be rendered following the same process as in Section III-B, with the color $c$ replaced by the normalized feature vector $\mathbf{f^{s}}$.
Training
We train the soft scale gate mapping $\mathcal{S}$ and Gaussian affinity feature $\mathbf{f}$ via scale-aware contrastive learning. Specifically, the correspondence distillation loss is given by:
$\mathcal{L}_{\mathrm{corr}}(s,p_{1},p_{2})=$ $\left(1-2\mathbbm{1}\left(\mathbf{V}(s,p_{1})\cdot\mathbf{V}(s,p_{2})>0\right)\right)$ (35)
$\max\left(\frac{\mathbf{F}^{s}(p_{1})\cdot\mathbf{F}^{s}(p_{2})}{\|\mathbf{F}^{s}(p_{1})\|\|\mathbf{F}^{s}(p_{2})\|}\right),$
@@
We consider two means of semantic analysis. For automatic decomposition of a dynamic scene, 4D Gaussians are clustered by their smoothed features (see colored clusters in Section VIII-G). This allows consistent 2D segmentation results to be rendered from any viewpoint at any timestamp. In addition, we enable 4D segmentation using 2D point prompts from a specific view. By extracting a 2D feature from the rendered feature map as input, we filter the 4D Gaussians according to the similarity between their features and the 2D query features. By integrating the scale mapping and scale-gated features, both segmentation algorithms support multi-granularity, controlled by the input scale.
VIII Experiments
VIII-A Experimental setup
VIII-A1 Datasets
Plenoptic Video dataset [4]
comprises six real-world scenes, each lasting ten seconds. For each scene, one view is reserved for testing while other views are used for training. To initialize the Guassians for this dataset, we utilize the colored point cloud generated by COLMAP from the first frame of each scene. The timestamps of each point are uniformly distributed across the scene’s duration. Following common practice, we downsample the resolution by 2×2$\times$2 × in training and evaluation.
D-NeRF dataset [9]
is a monocular video dataset comprising eight videos of synthetic scenes. Notably, at each time step, only a single image from a specific viewpoint is accessible. To assess model performance, we employ standard test views that originate from novel camera positions not encountered during the training process. The test views are taken within the time range covered by the training video. In this dataset, we utilize 100,000 randomly selected points, evenly distributed within the cubic volume defined by $[-1.2,1.2]^{3}$, and set their initial mean as the scene’s time duration.
VIII-A2 Implementation details
@@
To assess the versatility of our approach, we did not extensively fine-tune the training schedule across different datasets. By default, we conducted training with a total of 30,000 iterations, a batch size of 4, and halted densification at the midpoint of the schedule. We adopted the settings of [24] for hyperparameters such as loss weight, learning rate, and threshold. At the outset of training, we initialized both $q_{l}$ and $q_{r}$ as unit quaternions to establish identity rotations and set the initial time scaling to half of the scene’s duration. While the 4D Gaussian theoretically extends infinitely, we applied a Gaussian filter with marginal $p(t)<0.05$ when rendering the view at time $t$. For scenes in the Plenoptic Video dataset, we further initialized the Gaussian with 100,000 extra points distributed uniformly on the sphere encompassing the entire scene to fit the distant background that colmap failed to reconstruct and terminate its optimization after 10,000 iterations. Following the previous work, the LPIPS [94] in the Plenoptic Video dataset and the D-NeRF dataset are computed using AlexNet [95] and VGGNet [96] respectively.
VIII-B Results of dynamic novel view synthesis
TABLE I: Comparison with the state-of-the-art methods on the Plenoptic Video benchmark. 1: Only report the result on the scene flames salmon. 2: Only report SSIM instead of MS-SSIM like others. 3: Measured by ourselves using their official released code. 4: Results on Spinach, Beef, and Steak scenes.
Method PSNR $\uparrow$ DSSIM $\downarrow$ LPIPS $\downarrow$ FPS $\uparrow$
- Plenoptic Video (real, multi-view)
Neural Volumes [97]1 22.80 0.062 0.295 -
LLFF [98]1 23.24 0.076 0.235 -
DyNeRF [4]1 29.58 0.020 0.099 0.015
HexPlane [7] 31.70 0.014 0.075 0.563
K-Planes-explicit [6] 30.88 0.020 - 0.233
@@ VIII-B Results of dynamic novel view synthesis
StreamRF [99]1 29.58 - - 8.3
NeRFPlayer [10] 30.69 0.0352 0.111 0.045
HyperReel [8] 31.10 0.0372 0.096 2.00
4DGS-HexPlanes [42]4 31.02 0.030 0.150 36
@@
Results on the multi-view real scenes
Table I presents a quantitative evaluation on the Plenoptic Video dataset. Our approach not only significantly surpasses previous methods in terms of rendering quality but also achieves substantial speed improvements. Notably, it stands out as the sole method capable of real-time rendering while delivering high-quality dynamic novel view synthesis within this benchmark. To complement this quantitative assessment, we also offer qualitative comparisons on the “flame salmon” scene, as illustrated in Figure 3. The quality of synthesis in dynamic regions notably excels when compared to other methods. Several intricate details, including the black bars on the flame gun, the fine features of the right-hand fingers, and the texture of the salmon, are faithfully reconstructed, demonstrating the strength of our approach.
Results on the monocular synthetic videos
We also evaluate our approach on monocular dynamic scenes, a task known for its inherent complexities. Previous successful methods often rely on architectural priors to handle the underlying topology, but we refrain from introducing such assumptions when applying our 4D Gaussian model to monocular videos. Remarkably, our method surpasses all competing methods, as illustrated in Table II. This outcome underscores the ability of our 4D Gaussian model to efficiently exchange information across various time steps.
TABLE II: Qualitative comparison on monocular dynamic scenes. The results are averaged over all scenes in the D-NeRF dataset. 1: rendering at 800$\times$800, otherwise downsampled 2x by default.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
- D-NeRF (synthetic, monocular)
T-NeRF [9] 29.51 0.95 0.08
D-NeRF [9] 29.67 0.95 0.07
TiNeuVox [100] 32.67 0.97 0.04
HexPlanes [7] 31.04 0.97 0.04
K-Planes-explicit [6] 31.05 0.97 -
K-Planes-hybrid [6] 31.61 0.97 -
V4D [101] 33.72 0.98 0.02
4DGS-HexPlanes [42]1 33.30 0.98 0.03
4DGS (Ours) 34.09 0.98 0.02
@@
Ours (114 fps) DyNeRF (0.015 fps) [4] K-Planes (0.23 fps) [6] NeRFPlayer (0.045 fps) [10] HyperReel (2.00 fps) [8]
@@
Ground truth Neural Volumes [97] LLFF [98] HexPlane (0.56 fps) [7] MixVoxels (16.7 fps) [5]
@@
VIII-C Ablation and analysis
TABLE III: Ablation studies. We ablate our framework on two representative real scenes, flame salmon and cut roasted beef, which have volumetric effects, non-Lambertian surfaces, and different lighting conditions. “No-4DRot” denotes restricting the space and time independent of each other.

### v2

@@ 3.3 Training
Optimization In optimization, we only use the rendering loss as supervision. In most cases, combining the representation introduced above with the default training schedule as in [23] is sufficient to yield satisfactory results. However, for scenes with more dramatic changes, we observe issues such as temporal flickering and jitter. We consider that this may arise from suboptimal sampling techniques. Rather than adopting the prior regularization, we discover that batch sampling in time turns out to be effective, resulting in more seamless and visually pleasing appearance of dynamic contents.
Densification in spacetime In terms of density control, simply considering the magnitude of spatial position gradients is insufficient to assess under- and over-reconstruction over time. To address this, we incorporate the gradients of $\mu_{t}$ as an additional density control indicator. Furthermore, in regions prone to over-reconstruction, we employ joint spatial and temporal position sampling during Gaussian splitting.
4 Compact 4D Gaussian splatting
4.1 Analysis on memory bottleneck
4DGS can be interpreted as modeling the motion of dynamic scenes using piecewise-linear function, analogous to how 3D Gaussian model geometry. Given sufficient coverage of training views, the piecewise-linear function can offer a better fitting capability than those modeling the motion with trigonometric or polynomial functions, since enough piecewise-linear functions can also approximate these functions. However, this leads to a large number of primitives to model the motion of the same object, resulting in considerable redundancy and substantial storage demands.
To mitigate the storage bottleneck suffered by 4DGS, we consider two perspectives: (i) lowering the encoding cost for the attributes of each Gaussian (Section 4.2); (ii) decreasing the total number of primitives (Section 4.3), resulting in a compact version of 4DGS, dubbed 4DGSC.
4.2 Parameter compression
@@
4.3 Insignificant Gaussian removal
@@
5 Urban scene 4D Gaussian
We further extend 4DGS to urban scenes where the unbounded range of the scene itself and limited views of driving videos present unique challenges. To address these issues, we tailor a set of customizations to adapt the proposed representation to such scenes. First, we leverage LiDAR point clouds, typically available with corresponding timestamps aggregated across all training frames, to initialize the 4D position of 4D Gaussians. As the distance to the sky is much bigger than the scene’s scale, we employ a learnable cube map, $M_{\mathrm{sky}}$, to model its appearance, which can be queried by view direction $\mathbf{r}$. For each pixel, the cube map is used to consume the remaining transmittance:
@@ 5 Urban scene 4D Gaussian
$\mathcal{L}_{\mathrm{urban}}=$ $\lambda_{\mathrm{l1}}\mathcal{L}_{\mathrm{l1}}$ (30)
$+$ $\lambda_{\mathrm{d-ssim}}\mathcal{L}_{\mathrm{d-ssim}}+\lambda_{\mathrm{sky}}\mathcal{L}_{\mathrm{sky}}$
$+$ $\lambda_{\mathrm{lidar}}\mathcal{L}_{\mathrm{lidar}}+\lambda_{\mathrm{covt}}\mathcal{L}_{\mathrm{covt}}$
$+$ $\lambda_{\mathrm{sparse}}\mathcal{L}_{\mathrm{sparse}}+\lambda_{\mathrm{rigid}}\mathcal{L}_{\mathrm{rigid}},$
@@
6 Generative 4D Gaussian
In the preliminary version, we demonstrated that our proposed representation is capable of creating 4D content from monocular video via a reconstruction-based pipeline. However, effectively addressing such an inherently ill-posed problem in the more general case requires stronger priors as regularization. Existing methods [73, 77, 76, 79, 78] achieve this by incorporating generative models. As a fundamental representation, our approach is also well-suited to such approaches. For fair comparison, we use only anchored images [77] and score distill sampling (SDS) [65] as supervision, independent of 4D representation.
@@ 6 Generative 4D Gaussian
$\mathcal{L}=$ $\lambda_{\text{img}}\sum_{i=1}^{N}\|\hat{\mathcal{I}}_{i}^{\text{ref}}-\mathcal{I}_{i}^{\text{ref}}\|_{2}^{2}+\lambda_{\text{img}}\sum_{i=1}^{N}\sum_{j=1}^{M}\|\hat{\mathcal{I}}_{i,j}-\mathcal{I}_{i,j}\|_{2}^{2}$
@@
where $\lambda_{*}$ denotes the balancing weight, $\hat{\mathcal{I}}_{i}^{\text{ref}}$ and $\hat{\mathcal{I}}_{i,j}$ are the rendered images under fixed reference camera poses at timestamp $i$, and $\mathcal{I}^{\text{random}}_{i}$ is the rendered image under randomly sampled camera pose at timestamp $i$. We note rendered images $\hat{\mathcal{I}}_{i}^{\text{ref}}$, $\hat{\mathcal{I}}_{i,j}$ and $\mathcal{I}^{\text{random}}_{i}$ are 4D representation specific.
7 Segment anything in 4D Gaussian
@@
Scale-gated features We need to obtain the mask scale $s_{\mathbf{M}}$ for a 2D mask $\mathbf{M}$. To that end, $\mathbf{M}$ is first projected into 4D space to get the projected point cloud $\mathcal{P}$. Then $s_{\mathbf{M}}$ is calculated from $\mathcal{P}$ as:
@@
$2\sqrt{\text{std}(\mathcal{X}(\mathcal{P}))^{2}+\text{std}(\mathcal{Y}(\mathcal{P}))^{2}+\text{std}(\mathcal{Z}(\mathcal{P}))^{2}+\text{std}(\mathcal{T}(\mathcal{P}))^{2}},$
@@
The 2D feature $\mathbf{F}^{s}$ at a specific view could be rendered following the same process as in Section 3.2, with the color $c$ replaced by the normalized feature vector $\mathbf{f^{s}}$.
Training We train the soft scale gate mapping $\mathcal{S}$ and Gaussian affinity feature $\mathbf{f}$ via scale-aware contrastive learning. Specifically, the correspondence distillation loss is given by:
$\mathcal{L}_{\mathrm{corr}}(s,p_{1},p_{2})=$ $\left(1-2\mathbbm{1}\left(\mathbf{V}(s,p_{1})\cdot\mathbf{V}(s,p_{2})>0\right)\right)$
$\max\left(\frac{\mathbf{F}^{s}(p_{1})\cdot\mathbf{F}^{s}(p_{2})}{\|\mathbf{F}^{s}(p_{1})\|\|\mathbf{F}^{s}(p_{2})\|}\right),$ (34)
@@
Inference We consider two means of semantic analysis. For automatic decomposition of a dynamic scene, 4D Gaussians are clustered by their smoothed features (see colored clusters in Section 8.5). This allows consistent 2D segmentation results to be rendered from any viewpoint at any timestamp. In addition, we enable 4D segmentation using 2D point prompts from a specific view. By extracting a 2D feature from the rendered feature map as input, we filter the 4D Gaussians according to the similarity between their features and the 2D query features. By integrating the scale mapping and scale-gated features, both segmentation algorithms support multi-granularity, controlled by the input scale.
8 Experiments
8.1 4D from multi-view videos
8.1.1 Experimental setup
@@
Implementation details To assess the versatility of our approach, we did not extensively fine-tune the training schedule across different datasets. By default, we conducted training with a total of 30,000 iterations, a batch size of 4, and halted densification at the 15,000 iteration. We adopted the settings of [23] for hyperparameters such as loss weight, learning rate, and threshold. Before optimization, we initialized both $q_{l}$ and $q_{r}$ as unit quaternions to construct identity rotation matrix and set the initial time scaling to half of the scene’s duration. While the 4D Gaussian theoretically extends infinitely, we applied a Gaussian filter with marginal $p(t)<0.05$ when rendering the view at time $t$. The LPIPS [98] are computed using AlexNet [99].
Plenoptic Video dataset [4] contains six real-world scenes, each lasting ten seconds. For each scene, one view is reserved for testing while the other views are used for training. To initialize the Gaussians for this dataset, we utilize the colored point cloud generated by COLMAP from the first frame of each scene. The timestamps of each point are uniformly distributed across the scene’s duration. Following common practice, we downsample the resolution by $2\times$ in training and evaluation. Since the colmap may fail to reconstruct the distant background out of windows in scenes flames salmon and coffee martini. We simply adopt an environment sphere to model this part and terminate its optimization after 10,000 iterations.
Technicolor dataset [100] comprises light field videos of 12 scenes captured by a $4\times 4$ camera rig. Following the evaluation protocol in previous arts [8, 101, 102], we evaluate our method on 5 scenes (Birthday, Fabien, Painter, Theater, Trains) holding out the camera at the second row and the second column for testing. We retain the spatial resolution of $2560\times 1920$ for training and evaluation. For a fair comparison, we initialized our 4D Gaussian from the same sparse sfm points as [101]. Since there is negligible risk of floaters in some scenes (e.g., Painter), we disable the periodic reset of opacity in these scenes to achieve better performance. The SSIM is evaluated using sckit-image library function with $datarange=1.0$.
8.1.2 Results
Table 1: Comparison with the state-of-the-art methods on the Plenoptic Video benchmark.
[…]
- Plenoptic Video
Neural Volumes [103] [^1] Only report the result on the scene flames salmon. 22.80 0.062 0.295 -
LLFF [104] [^1] Only report the result on the scene flames salmon. 23.24 0.076 0.235 -
DyNeRF [4] [^1] Only report the result on the scene flames salmon. 29.58 0.020 0.099 0.015
HexPlane [7] 31.70 0.014 0.075 0.56333Measured by ourselves using their officially released code.
K-Planes-explicit [6] 30.88 0.020 - 0.23333Measured by ourselves using their officially released code.
@@ 8.1.2 Results
StreamRF [105] [^1] Only report the result on the scene flames salmon. 29.58 - - 8.3
NeRFPlayer [10] 30.69 0.035222Only report SSIM instead of MS-SSIM like others. 0.111 0.045
HyperReel [8] 31.10 0.037222Only report SSIM instead of MS-SSIM like others. 0.096 2.00
4DGaussian [41] [^4] Results on Spinach, Beef, and Steak scenes. 31.02 0.030 0.150 36
4D-Rotor-Gaussian [44] 31.62 0.030 0.140 277
@@
Ours DyNeRF [4] K-Planes [6] NeRFPlayer [10] HyperReel [8]
(114 fps) (0.015 fps) (0.23 fps) (0.045 fps) (2.00 fps)
@@
Ground truth Neural Volumes [103] LLFF [104] HexPlane [7] MixVoxels [5]
- - - (0.56 fps) (16.7 fps)
@@
Evaluation on Plenoptic Video dataset [4]. Table 1 presents a quantitative evaluation on the Plenoptic Video dataset. Our approach not only significantly surpasses previous methods in terms of rendering quality but also achieves substantial speed improvements. Notably, it stands out as the sole method capable of real-time rendering while delivering high-quality dynamic novel view synthesis within this benchmark. To complement this quantitative assessment, we also offer qualitative comparisons on the “flame salmon” scene, as illustrated in Figure 3. The quality of synthesis in dynamic regions notably excels when compared to other methods. Several intricate details, including the black bars on the flame gun, the fine features of the right-hand fingers, and the texture of the salmon, are faithfully reconstructed, demonstrating the strength of our approach.
Table 2: Quantitative comparison on the Technicolor dataset.
[…]
- Technicolor
DyNeRF [4] [^1] Reported from [102]. 31.80 - 0.142 0.6 MB [^2] Reported from [101].
HyperReel [8] [^1] Reported from [102]. 32.73 0.906 0.109 1.2 MB [^2] Reported from [101].
4DGaussian [41] [^1] Reported from [102]. 30.79 0.843 0.178 1.0 MB [^3] Reported from [106].
Ex4DGS [102] [^1] Reported from [102]. 33.62 0.916 0.088 1.5 MB [^3] Reported from [106].
STG [101] [^2] Reported from [101]. 33.56 0.920 0.084 1.1 MB [^2] Reported from [101].
4DGS 34.03 0.921 0.084 37.6 MB
4DGSC [^4] Compact variant of our 4DGS. 33.75 0.913 0.865 1.6 MB
Evaluation on Technicolor dataset. To validate the applicability to diverse real-world scenes, we compare our approach with previous methods on the Technicolor dataset. The qualitative results in Figure 4 demonstrate that our method has superior capability for capturing subtle details in the dynamic regions. The quantitative results in Table 2 suggest that our method achieves the highest performance across quality-related metrics. While there seems to be a storage bottleneck compared to other baselines since the vanilla 4DGS focuses on boosting the quality, this can be easily addressed. As shown in the table, when equipped with the compression strategy introduced in Section 4, the method (4DGSC) achieves a significant improvement in compactness without significant compromise in quality.
Refer to caption Refer to caption Refer to caption Refer to caption
GT HyperReel [8] STG [101] Ours
Figure 4: Qualitative comparison on the Technicolor dataset.
8.1.3 Ablation and analysis
Table 3: Ablation studies. We ablate our framework on two representative real scenes, flame salmon and cut roasted beef, which have volumetric effects, non-Lambertian surfaces, and different lighting conditions. “No-4DRot” denotes restricting the space and time independent of each other.